OpenStudy (anonymous):

Simplify y= a ∙ (1 - 0.09) ^x

OpenStudy (anonymous):

i have the answer

OpenStudy (anonymous):

(3xy2 + 100) • (3xy2 - 100) ——————————————————————————— 100

OpenStudy (anonymous):

Reformatting the input : Changes made to your input should not affect the solution: (1): "0.09" was replaced by "(09/100)". Step by step solution : Step 1 : Raise y to the 4th power Exponentiation : Equation at the end of step 1 : 9 ((——— • (x2)) • y4) - 100 100 Step 2 : Raise x to the 2nd power Exponentiation : Equation at the end of step 2 : 9 ((——— • x2) • y4) - 100 100 Step 3 : 9x2y4 Simplify ————— - 100 100 Rewriting the whole as an Equivalent Fraction : 3.1 Subtracting a whole from a fraction Rewrite the whole as a fraction using 100 as the denominator : 100 100 • 100 100 = ——— = ————————— 1 100 Equivalent fraction : The fraction thus generated looks different but has the same value as the whole Common denominator : The equivalent fraction and the other fraction involved in the calculation share the same denominator Adding fractions that have a common denominator : 3.2 Adding up the two equivalent fractions Add the two equivalent fractions which now have a common denominator Combine the numerators together, put the sum or difference over the common denominator then reduce to lowest terms if possible: 9x2y4 - (100 • 100) 9x2y4 - 10000 ——————————————————— = ————————————— 100 100 Trying to factor as a Difference of Squares : 3.3 Factoring: 9x2y4 - 10000 Theory : A difference of two perfect squares, A2 - B2 can be factored into (A+B) • (A-B) Proof : (A+B) • (A-B) = A2 - AB + BA - B2 = A2 - AB + AB - B2 = A2 - B2 Note : AB = BA is the commutative property of multiplication. Note : - AB + AB equals zero and is therefore eliminated from the expression. Check : 9 is the square of 3 Check : 10000 is the square of 100 Check : x2 is the square of x1 Check : y4 is the square of y2 Factorization is : (3xy2 + 100) • (3xy2 - 100) Trying to factor as a Difference of Squares : 3.4 Factoring: 3xy2 - 100 Check : 3 is not a square !! Ruling : Binomial can not be factored as the difference of two perfect squares Final result : (3xy2 + 100) • (3xy2 - 100) ——————————————————————————— 100

OpenStudy (anonymous):

its an exponential function i dont think ur correct

OpenStudy (anonymous):

(9/100)x2y4-100 Final result : (3xyover2 + 100) • (3xyover2 - 100) ——————————————————————————— 100

OpenStudy (anonymous):

ok let me write it for you

OpenStudy (anonymous):

understand now @babygirlang66666

OpenStudy (anonymous):

?

OpenStudy (anonymous):

thats not right it should be a deciaml

OpenStudy (anonymous):

its an equation honey, you have to solve it.

OpenStudy (anonymous):

well im not good at this, thats why im on here.. i have this last question for this project and i need to turn it in

OpenStudy (anonymous):

so can u just simplify it for me

OpenStudy (anonymous):

you have to understand that here we dont give you the answers we explain how to do it. OK? we dont do your homework honey, we help you.

OpenStudy (anonymous):

i know... but idk how to do it

OpenStudy (anonymous):

so if you dont know what 3xy2 + 100 is, pull out a calculator dear.

OpenStudy (anonymous):

is this a multiple choice?

OpenStudy (anonymous):

no ill show u the full question hold on

OpenStudy (anonymous):

b) You are in charge. By what percentage will you tell Greece to cut their spending? What is the decay factor? 9% y= a ∙ (1 - 0.09) ^x y= a ∙ (1 - b=

OpenStudy (anonymous):

it is 174.24

OpenStudy (anonymous):

to the 1st question

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem, can you click the best response button next to my comment?